A sled is pulled from rest by a net force of 2.8 N [forward]. The sled moves 7.4 m, reaching a velocity of 3.9  m/s [forward]. Determine the mass of the sled.


1
Expert's answer
2022-04-19T12:06:07-0400

The energy-work theorem says

"\frac{mv^2}{2}=Fd"

Hence, the mass of the object

"m=\frac{2Fd}{v^2}"

"m=\frac{2*2.8*7.4}{3.9^2}=2.72\:\rm kg"